Please help me on this one. I get confused with Word Problems!!! Please help me :) Thanks so much!!!

J.Bond and E.Blofeld are driving speedboats in opposite directions after a failed attempt by the latter to blow up the world with a nuclear bomb. Mr.Bonds boat is going 25 MPH faster than Mr.Blofelds. After 3 hours they are 500 miles apart. How fast was each of them going?

Answer by rfer(16322)

x+x+25=500/3
2x+25=167
2x=142
x=71 mph
x+25=96 mph